# Robust targeted exploration for systems with non-stochastic disturbances

### v2

@@ Abstract
In this paper, we introduce a novel targeted exploration strategy designed specifically for uncertain linear time-invariant systems with energy-bounded disturbances, i.e., without making any assumptions on the distribution of the disturbances. We use classical results characterizing the set of non-falsified parameters consistent with energy-bounded disturbances. We derive a semidefinite program which computes an exploration strategy that guarantees a desired accuracy of the parameter estimate. This design is based on sufficient conditions on the spectral content of the exploration data that robustly account for initial parametric uncertainty. Finally, we highlight the applicability of the exploration strategy through a numerical example involving a nonlinear system.
@@ 1 Introduction
Designing reliable controllers for unknown dynamical systems requires accurate knowledge of the model parameters, which can be obtained from data [gevers2005identification]. The accuracy of the parameters significantly depends on the quality of the data used for system identification. Informative data can be strategically obtained from an experiment through the process of targeted exploration or optimal experiment design [pronzato2008optimal, gevers1986optimal]. Specifically, targeted exploration inputs are tailored to reduce model uncertainty, thereby ensuring the attainment of a desired accuracy in the identified model [jansson2005input, bombois2021robust], or the feasibility of robust control design  [barenthin2008identification, umenberger2019robust, ferizbegovic2019learning, venkatasubramanian2020robust, venkatasubramanian2023sequential]. In this paper, we develop a targeted exploration strategy for uncertain linear systems subject to energy-bounded disturbances.
Targeted exploration is typically studied in a stochastic setup, for example, with independent and identically distributed (i.i.d.) disturbances with zero mean. In this case, one can construct a data-based confidence ellipsoid for the parameters [ljung1999system] that can be approximately predicted and optimized. This classical asymptotic result has been utilized to design targeted exploration strategies for dual control methods [barenthin2008identification, umenberger2019robust, iannelli2019structured, venkatasubramanian2020robust, venkatasubramanian2023sequential]. In [dean2017sample], tight confidence regions for the parameters are constructed, assuming the data is independent, and hence it is not applicable to correlated time-series data from a single trajectory. Nevertheless, this bound has been utilized in the design of targeted exploration for dual control in [ferizbegovic2019learning].
@@
A common feature among all of the discussed targeted exploration approaches is that they consider linear systems subject to i.i.d. zero mean stochastic disturbances. However, real-world systems often exhibit nonlinear behaviour or unmodeled dynamics, i.e., errors in the assumed model structure, which introduce additional deterministic model mismatch and cannot be explained by independent stochastic noise [sarker2023accurate]. Instead, such disturbances can be modeled as bounded disturbances, assuming that the disturbances belong to a known bounded set, e.g., point-wise bounds or energy bounds. To quantify models under such disturbances, set-membership estimation methods have been developed, and various identification results in this direction can be found in [fogel1979system, milanese1989estimation, sarker2023accurate, bisoffi2021trade, shakouri2024identification]. Although these methods have recently gained popularity for data-driven robust control design [van2020noisy, berberich2022combining], a principled method for optimal experiment design in the presence of bounded disturbances is still lacking.
@@
In this paper, we design a targeted exploration strategy that ensures a desired error bound on the estimated parameters by utilizing a data-dependent uncertainty bound based on energy-bounded disturbances [fogel1979system]. In contrast to existing optimal experiment design methods, which often rely on stochastic disturbance assumptions [umenberger2019robust, ferizbegovic2019learning, venkatasubramanian2023sequential], our approach provides robustness against non-stochastic, adversarial disturbances and is applicable to systems with bounded nonlinearities. We consider multi-sine exploration inputs of specific frequencies and optimized amplitudes to explicitly shape and reduce uncertainty in a targeted manner. As one of our main contributions, we derive a sufficient condition on the spectral content of the exploration data that asymptotically guarantees a desired error-bound on the parameters estimated through exploration. We utilize the sufficient condition on the spectral content to derive LMIs for exploration which ensure the desired error bound on the parameters. This approach gives rise to a targeted exploration design with minimal input energy based on a semidefinite program (SDP). While existing work has also derived SDPs for targeted exploration design [venkatasubramanian2023sequential], the non-stochastic identification results considered in the proposed work yield structurally different requirements on excitation, which require new worst-case robust guarantees rather than high-probability guarantees. To this end, we provide the setting and define the exploration goal in Section 2, and the preliminaries regarding data-dependent uncertainty bounds in Section 3. In Section 4, we derive the exploration strategy by leveraging a sufficient condition on the time-series exploration data, as proposed in our previous work [venkatasubramanian2023towards]. Furthermore, we robustly account for parametric uncertainty and error due to disturbances with suitable bounds in Section 4, unlike [venkatasubramanian2023towards]. Finally, in Section 5, we provide a numerical example to highlight the applicability of the proposed exploration strategy to nonlinear systems.
@@ 2 Problem statement
Notation: The transpose of a matrix $A\in\mathbb{R}^{n\times m}$ is denoted by $A^{\top}$. The conjugate transpose of a matrix $A\in\mathbb{C}^{n\times m}$ is denoted by $A^{\mathsf{H}}$. The positive definiteness and positive semi-definiteness of a matrix $A\in\mathbb{C}^{n\times n}$ is denoted by $A=A^{\mathsf{H}}\succ 0$ and $A=A^{\mathsf{H}}\succeq 0$, respectively. The operator $\mathrm{vec}(A)$ stacks the columns of $A$ to form a vector. The operator $\textnormal{diag}(A_{1},\dots,A_{n})$ creates a block diagonal matrix by aligning the matrices $A_{1},\dots,A_{n}$ along the diagonal starting with $A_{1}$ in the upper left corner. The Kronecker product is denoted by $\otimes$. The Euclidean norm and the weighted Euclidean norm for a vector $x\in\mathbb{R}^{n}$ and a matrix $P\succ 0$ are denoted by $\|x\|=\sqrt{x^{\top}x}$ and $\|x\|_{P}=\sqrt{x^{\top}Px}$, respectively. The largest singular value of a matrix $A\in\mathbb{C}^{m\times n}$ is denoted by $\|A\|$. Furthermore, given a matrix $M\succeq 0$, $\|A\|_{M}=\|M^{1/2}A\|$ where $M^{1/2}$ is the symmetric square root matrix of $M$. The identity matrix of size $n$ is denoted by $I_{n}$. A vector of ones of size $n$ is denoted by $\mathbf{1}_{n}\in\mathbb{R}^{n\times 1}$.
Amplitude of a spectral line [venkatasubramanian2023sequential, Def. 2]: Given a sequence $\{\phi_{k}\}_{k=0}^{T-1}$, the amplitude of the spectral line of the sequence at a frequency $\omega\in\Omega_{T}:=\{0,1/T,\dots,(T-1)/T\}$ is given by
@@ 2.1 Setting
where $x_{k}\in\mathbb{R}^{n_{\mathrm{x}}}$ is the state, $u_{k}\in\mathbb{R}^{n_{\mathrm{u}}}$ is the control input, and $w_{k}\in\mathbb{R}^{n_{\mathrm{x}}}$ is the disturbance. In our setting, the true system parameters $A_{\mathrm{tr}}$, $B_{\mathrm{tr}}$ are uncertain. Hence, it is necessary to collect informative data from an optimal experiment for a fixed $T\in\mathbb{N}$ time steps to enhance the accuracy of the parameters. It is assumed that the state can be measured, the initial state is at the origin, i.e., $x_{0}=0$, and the disturbances are energy-bounded.
@@
The disturbances $w$ are energy-bounded, i.e., there exists a known constant $\gamma_{\mathrm{w}}>0$ such that
@@
This assumption on the disturbances allows the system to exhibit nonlinear behaviour, as long as the nonlinearities can be bounded in energy (as discussed in Section 5). Importantly, this assumption represents the main difference to existing experiment design methods which are typically restricted to i.i.d. Gaussian disturbances, e.g., [umenberger2019robust, ferizbegovic2019learning, venkatasubramanian2023sequential].
@@ 2.2 Exploration strategy
where $T$ is the exploration time and $\bar{u}(\omega_{i})\in\mathbb{R}^{n_{\mathrm{u}}}$ are the amplitudes of the sinusoidal inputs at $L\in\mathbb{N}$ distinct selected frequencies $\omega_{i}\in\Omega_{T}$ with $n_{\mathrm{x}}+n_{\mathrm{u}}\leq L\leq T$. In practice, frequencies may be selected based on prior information [mehra1974optimal]. Since the input signal is deterministic and sinusoidal, the amplitude of the spectral line of the sequence $\{u_{k}\}_{k=0}^{T-1}$ at frequency $\omega_{i}$ is $\bar{u}(\omega_{i})$. Denote $U_{\mathrm{e}}=\mathrm{diag}(\bar{u}(\omega_{1}),\dots,\bar{u}(\omega_{L}))\in\mathbb{R}^{Ln_{\mathrm{u}}\times L}$. The exploration input is computed such that it excites the system sufficiently with minimal input energy, based on the initial parameter estimates. Bounding the input energy by a constant $\gamma^{2}_{\mathrm{e}}$ can be equivalently written as $\sum_{i=1}^{L}\|\bar{u}(\omega_{i})\|^{2}=\mathbf{1}_{L}^{\top}U_{\mathrm{e}}^{\top}U_{\mathrm{e}}\mathbf{1}_{L}\preceq\gamma_{\mathrm{e}}^{2}$ where $\mathbf{1}_{L}\in\mathbb{R}^{L\times 1}$ is a vector of ones, and the bound $\gamma_{\mathrm{e}}\geq 0$ is desired to be small. Using the Schur complement, this criterion is equivalent to
@@
We require $A_{\mathrm{tr}}$ to be Schur stable since we consider only open-loop inputs in our exploration strategy (6). Assumption 4 could be relaxed if an exploration input of the form in (6) with an additional linear feedback, i.e., $v_{k}=u_{k}+Kx_{k}$, is utilized which ensures robust stability for all $\theta\in\mathbf{\Theta}_{0}$ (5).
In order to achieve the exploration goal, the amplitudes of the sinusoidal exploration inputs need to be optimized such that by applying the exploration inputs, the obtained estimate satisfies the desired uncertainty bound (4).
@@ 3 Preliminaries on data-driven uncertainty quantification
In this section, we discuss a data-dependent uncertainty bound on the parameter estimates in the presence of energy-bounded noise [fogel1979system]. Given observed data $\mathcal{D}_{T+1}=\{x_{k},u_{k}\}_{k=0}^{T}$ of length $T+1$, the objective is to quantify the uncertainty associated with the unknown parameters $\theta_{\mathrm{tr}}$. Henceforth, we denote $\phi_{k}=[x_{k}^{\top}\,u_{k}^{\top}]^{\top}\in\mathbb{R}^{n_{\phi}}$ where $n_{\phi}=n_{\mathrm{x}}+n_{\mathrm{u}}$. The system (2) can be re-written in terms of parameter $\theta_{\mathrm{tr}}=\textnormal{vec}([A_{\mathrm{tr}},B_{\mathrm{tr}}])$ as
$x_{k+1}=(\phi_{k}^{\top}\otimes I_{n_{\mathrm{x}}})\theta_{\mathrm{tr}}+w_{k}.$ (8)
@@
$\Phi=[\phi_{0},\dots,\phi_{T-1}]\in\mathbb{R}^{n_{\phi}\times T}$ (9)
and
$X^{\top}=[x_{1}^{\top},\dots,x_{T}^{\top}]\in\mathbb{R}^{1\times Tn_{\mathrm{x}}}.$ (10)
We obtain the following expressions for the mean $\hat{\theta}_{T}~=~\textnormal{vec}([\hat{A}_{T},\hat{B}_{T}])$ and covariance $P$ of the parameters from the standard least squares formulation [ljung1999system, Section 1.3]:
$\hat{\theta}_{T}=P\sum_{k=0}^{T-1}(\phi_{k}^{\top}\otimes I_{n_{\mathrm{x}}})^{\top}x_{k+1}=P(\Phi\otimes I_{n_{\mathrm{x}}})X$ (11)
and
@@
[fogel1979system] Let Assumption 1 hold. Given data set $\mathcal{D}_{T+1}$, the set of non-falsified parameters $\theta$ is given by
@@
Given Assumption 1, the non-falsified set $\bm{\Theta}_{T}$ provides an exact characterization of the set of parameters explaining the data. Similar non-falsified sets for matrices are considered in [van2020noisy, berberich2022combining]. The ellipsoid (13), derived from energy-bounded constraints, is characterized by a vector $\hat{\theta}_{T}$ and a matrix $P$, which correspond to the mean and covariance of the least squares estimator for linear systems with Gaussian disturbances [umenberger2019robust, Prop. 2.1]. However, unlike the case of zero mean i.i.d. Gaussian disturbances, in the case of energy-bounded disturbances the scaling $G$ (14) of the bounding ellipsoid is also data-dependent. This dependence has significant consequences on optimal experiment design, as will be discussed in the next section. Furthermore, for $\gamma_{\mathrm{w}}\propto T$, $G$ scales at most linearly with $T$ as $T\to\infty$ since $G\leq\gamma_{\mathrm{w}}$ [fogel1979system, Lemma 4]. In contrast, the scaling of the confidence ellipsoid in the Gaussian case does not depend on $T$ [ljung1999system, umenberger2019robust]. Since $P^{-1}$ increases linearly with $T$, the size of the confidence ellipsoid in the Gaussian case reduces with $T$ [ljung1999system, umenberger2019robust]. However, in the considered case of energy-bounded disturbances, the size of the non-falsified set $\mathbf{\Theta}_{T}$, in general, does not decrease as $T\to\infty$.
@@ 4 Targeted Exploration
In this section, we propose a targeted exploration strategy based on the data-dependent uncertainty bound provided in Lemma 6. The exploration strategy builds upon sufficient conditions on the exploration data outlined in Section 4.1. In particular, we derive bounds on the exploration data using the spectral information of the exploration inputs in Section 4.2. Since these bounds depend on the uncertain model parameters, bounds on the effect of model uncertainty are derived in Section 4.3. We utilize these bounds to derive sufficient conditions on the spectral information of the exploration inputs. However, since the derived sufficient conditions are non-convex in the decision variables, a convex relaxation procedure is carried out in Section 4.5. Finally, in Section 4.6, the exploration problem is reduced to a set of LMIs that provide us exploration inputs that ensure the exploration goal.

### v3

@@ Abstract
We propose a novel targeted exploration strategy designed specifically for uncertain linear time-invariant systems with energy-bounded disturbances, i.e., without any assumptions on the distribution of the disturbances. We use classical results characterizing the set of non-falsified parameters consistent with energy-bounded disturbances. We derive a semidefinite program which computes an exploration strategy that guarantees a desired accuracy of the parameter estimate. This design is based on sufficient conditions on the spectral content of the exploration data that robustly account for initial parametric uncertainty. Finally, we highlight the applicability of the exploration strategy through a numerical example involving a nonlinear system.
@@ 1 Introduction
Designing reliable controllers for unknown dynamical systems requires accurate knowledge of the model parameters, which can be obtained from data [11]. The accuracy of the parameters significantly depends on the quality of the data used for system identification. Informative data can be strategically obtained from an experiment through the process of targeted exploration or optimal experiment design [19, 10]. Specifically, targeted exploration inputs are tailored to reduce model uncertainty, thereby ensuring the attainment of a desired accuracy in the identified model [14, 4], or the feasibility of robust control design  [1, 23, 8, 26]. In this paper, we develop a targeted exploration strategy for uncertain linear systems subject to energy-bounded disturbances.
@@
Targeted exploration is typically studied in a stochastic setup, for example, with independent and identically distributed (i.i.d.) disturbances with zero mean. In this case, one can construct a data-based confidence ellipsoid for the parameters [15] that can be approximately predicted and optimized. This classical asymptotic result has been utilized to design targeted exploration strategies for dual control methods [1, 23, 13, 26]. In [7], tight confidence regions for the parameters are constructed, assuming the data is independent, and hence it is not applicable to correlated time-series data from a single trajectory. Nevertheless, this bound has been utilized in the design of targeted exploration for dual control in [8]. A common feature among all of the discussed targeted exploration approaches is that they consider linear systems subject to i.i.d. zero mean stochastic disturbances. However, real-world systems often exhibit nonlinear behaviour or unmodeled dynamics which introduce additional deterministic model mismatch and cannot be explained by independent stochastic noise [20]. Instead, such effects can be modeled as bounded disturbances, assuming that the disturbances belong to a known bounded set, e.g., point-wise bounds or energy bounds. To quantify models under such disturbances, set-membership estimation methods have been developed, and various identification results in this direction can be found in [9, 18, 20, 3, 22]. Although these methods have recently gained popularity for data-driven robust control design [24, 2], a principled method for optimal experiment design in the presence of bounded disturbances is still lacking.
@@
In this paper, we design a targeted exploration strategy that ensures a desired error bound on the estimated parameters by utilizing a data-dependent uncertainty bound based on energy-bounded disturbances [9]. Unlike existing experiment design methods, which often rely on stochastic disturbance assumptions [23, 8, 26], our approach provides robustness against non-stochastic, adversarial disturbances and is applicable to systems with bounded nonlinearities. We consider multi-sine exploration inputs of specific frequencies and optimized amplitudes to explicitly shape and reduce uncertainty in a targeted manner. As a main contribution, we derive sufficient conditions on the spectral content of the exploration data that asymptotically guarantees a desired error-bound on the parameters estimated through exploration. We utilize these sufficient conditions to derive LMIs for exploration which ensure the desired error bound on the parameters. This approach gives rise to a targeted exploration design with minimal input energy based on a semidefinite program (SDP). Unlike existing methods for stochastic disturbances [26, 23, 8], the non-stochastic identification results considered in this work yield structurally different requirements on excitation, which require new worst-case robust guarantees rather than high-probability guarantees. To this end, we formulate the problem in Section 2, and review data-dependent uncertainty bounds in Section 3. In Section 4, we derive the exploration strategy by leveraging a sufficient conditions on time-series exploration data, as proposed in our previous work [28], while additionally robustly accounting for parametric uncertainty and error due to disturbances with suitable bounds. In Section 5, we provide a numerical example to highlight the applicability of the proposed exploration strategy to nonlinear systems. We remark that the proofs of Lemma 9, Propositions 11 and 12, and Theorem 13 are not included in this brief paper, but can be found in the technical report [27].
@@ 2 Problem statement
Notation: For a matrix $\in\mathbb{R}^{n\times m}, AA^{\top}$ denotes the transpose. For a matrix $A\in\mathbb{C}^{n\times m}$, $A^{\mathsf{H}}$ denotes the conjugate transpose. The positive (semi-)definiteness of a matrix $A\in\mathbb{C}^{n\times n}$ is denoted by ($A=A^{\mathsf{H}}\succeq 0$) $A=A^{\mathsf{H}}\succ 0$. The operator $\mathrm{vec}(A)$ stacks the columns of $A$ to form a vector. The operator $\textnormal{diag}(A_{1},\dots,A_{n})$ creates a block diagonal matrix by aligning the matrices $A_{1},\dots,A_{n}$ along the diagonal with $A_{1}$ in the upper left corner. The Kronecker product is denoted by $\otimes$. The Euclidean norm and the weighted Euclidean norm for a vector $x\in\mathbb{R}^{n}$ and a matrix $P\succ 0$ are denoted by $\|x\|=\sqrt{x^{\top}x}$ and $\|x\|_{P}=\sqrt{x^{\top}Px}$, respectively. The largest singular value of a matrix $A\in\mathbb{C}^{m\times n}$ is $\|A\|$. Furthermore, given a matrix $M\succeq 0$, $\|A\|_{M}=\|M^{1/2}A\|$ where $M^{1/2}$ is the symmetric square root matrix of $M$. The identity matrix of size $n$ is denoted by $I_{n}$. A vector of ones of size $n$ is denoted by $\mathbf{1}_{n}\in\mathbb{R}^{n\times 1}$.
Amplitude of a spectral line [26, Def. 2]: The amplitude of the spectral line of a sequence $\{\phi_{k}\}_{k=0}^{T-1}$ at a frequency $\omega\in\Omega_{T}:=\{0,1/T,\dots,(T-1)/T\}$ is given by
@@ 2.1 Setting
where $x_{k}\in\mathbb{R}^{n_{\mathrm{x}}}$ is the state, $u_{k}\in\mathbb{R}^{n_{\mathrm{u}}}$ is the control input, and $w_{k}\in\mathbb{R}^{n_{\mathrm{x}}}$ is the disturbance. In our setting, the true system parameters $A_{\mathrm{tr}}$, $B_{\mathrm{tr}}$ are uncertain. Hence, it is necessary to collect informative data from an optimal experiment for a fixed $T\in\mathbb{N}$ time steps to improve parameter accuracy. To simplify the exposition, we assume that the initial state is at the origin, i.e., $x_{0}=0$. Furthermore, we assume that the state can be measured and the disturbances are energy-bounded.
@@
The disturbances $w$ are energy-bounded; there exists a known constant $\gamma_{\mathrm{w}}>0$ such that
@@
This assumption allows the system to exhibit nonlinear behaviour, as long as the nonlinearities can be bounded in energy (as discussed in Section 5). Importantly, this assumption distinguishes our approach from existing experiment design methods which are typically restricted to i.i.d. Gaussian disturbances, e.g., [23, 8, 26].
@@ 2.2 Exploration strategy
where $T$ is the exploration time and $\bar{u}(\omega_{i})\in\mathbb{R}^{n_{\mathrm{u}}}$ are the amplitudes of the sinusoidal inputs at $L\in\mathbb{N}$ distinct selected frequencies $\omega_{i}\in\Omega_{T}$ with $n_{\mathrm{x}}+n_{\mathrm{u}}\leq L\leq T$. In practice, frequencies may be selected based on prior information [17]. Since the input signal is deterministic and sinusoidal, the amplitude of the spectral line of the sequence $\{u_{k}\}_{k=0}^{T-1}$ at frequency $\omega_{i}$ is $\bar{u}(\omega_{i})$. Denote $U_{\mathrm{e}}=\mathrm{diag}(\bar{u}(\omega_{1}),\dots,\bar{u}(\omega_{L}))\in\mathbb{R}^{Ln_{\mathrm{u}}\times L}$. The exploration input is computed such that it excites the system sufficiently with minimal input energy, based on the initial parameter estimates. Bounding the input energy by a constant $\gamma^{2}_{\mathrm{e}}$ is equivalent to $\sum_{i=1}^{L}\|\bar{u}(\omega_{i})\|^{2}=\mathbf{1}_{L}^{\top}U_{\mathrm{e}}^{\top}U_{\mathrm{e}}\mathbf{1}_{L}\preceq\gamma_{\mathrm{e}}^{2}$ where $\mathbf{1}_{L}\in\mathbb{R}^{L\times 1}$ is a vector of ones, and the bound $\gamma_{\mathrm{e}}\geq 0$ is desired to be small. Using the Schur complement, this criterion is equivalent to
@@
$A_{\mathrm{tr}}$ is required to be Schur stable since we consider only open-loop inputs in our exploration strategy (6). Assumption 4 could be relaxed by augmenting the inputs in (6) with an additional linear feedback, i.e., $v_{k}=u_{k}+Kx_{k}$, ensuring robust stability for all $\theta\in\mathbf{\Theta}_{0}$ (5).
To achieve the exploration goal, the amplitudes of the sinusoidal exploration inputs need to be optimized such that by applying the exploration inputs, the obtained estimate satisfies the desired uncertainty bound (4).
@@ 3 Preliminaries on data-driven uncertainty quantification
In this section, we discuss a data-dependent uncertainty bound on the parameter estimates in the presence of energy-bounded noise [9]. Given observed data $\mathcal{D}_{T+1}=\{x_{k},u_{k}\}_{k=0}^{T}$ of length $T+1$, the objective is to quantify the uncertainty associated with the unknown parameters $\theta_{\mathrm{tr}}$. Henceforth, we denote $\phi_{k}=[x_{k}^{\top}\,u_{k}^{\top}]^{\top}\in\mathbb{R}^{n_{\phi}}$ where $n_{\phi}=n_{\mathrm{x}}+n_{\mathrm{u}}$. The system (2) can be re-written in terms of parameter $\theta_{\mathrm{tr}}=\textnormal{vec}([A_{\mathrm{tr}},B_{\mathrm{tr}}])$ as
$x_{k+1}=(\phi_{k}^{\top}\otimes I_{n_{\mathrm{x}}})\theta_{\mathrm{tr}}+w_{k}.$
@@
$\Phi$ $=[\phi_{0},\dots,\phi_{T-1}]\in\mathbb{R}^{n_{\phi}\times T},$
$X^{\top}$ $=[x_{1}^{\top},\dots,x_{T}^{\top}]\in\mathbb{R}^{1\times Tn_{\mathrm{x}}}.$
We obtain the following expressions for the mean $\hat{\theta}_{T}~=~\textnormal{vec}([\hat{A}_{T},\hat{B}_{T}])$ and covariance $P$ of the parameters from the standard least squares formulation [15, Section 1.3]:
$\hat{\theta}_{T}$ $=P\sum_{k=0}^{T-1}(\phi_{k}^{\top}\otimes I_{n_{\mathrm{x}}})^{\top}x_{k+1}=P(\Phi\otimes I_{n_{\mathrm{x}}})X,$ (8)
@@
[9] Let Assumption 1 hold. Given data set $\mathcal{D}_{T+1}$, the set of non-falsified parameters $\theta$ is given by
@@ 4 Targeted Exploration
In this section, we propose a targeted exploration strategy based on the data-dependent uncertainty bound in Lemma 6. The exploration strategy builds upon sufficient conditions on the exploration data outlined in Section 4.1. In particular, we derive bounds on the exploration data using the spectral information of the exploration inputs in Section 4.2. Since these bounds depend on the uncertain model parameters, bounds on the effect of model uncertainty are derived in Section 4.3. We utilize these bounds to derive sufficient conditions on the spectral information of the exploration inputs in Section 4.4. However, since the derived sufficient conditions are non-convex in the decision variables, a convex relaxation procedure is carried out in Section 4.5. Finally, in Section 4.6, the exploration problem is reduced to a set of LMIs that provide us exploration inputs that ensure the exploration goal.